# Estimates for Dirichlet Eigenvalues of the Schrödinger operator with the Kronig-Penney Model

### v1

@@ Abstract
In this paper, first, we improve the asymptotic formulas obtained in [13] and obtain sharp asymptotic formulas explicitly expressed by the potential. For the potentials of bounded variation, we obtain asymptotic formulas in which the first and second terms are explicitly determined and separated from the error terms. In addition, we illustrate these formulas for the Kronig-Penney potential. We then provide estimates for the small Dirichlet eigenvalues of the one-dimensional Schrödinger operator in the Kronig-Penney model. Using Rouché’s theorem, we derive several useful equations from certain iteration formulas for computing these Dirichlet eigenvalues, and we estimate the eigenvalues numerically. Moreover, we present error estimates and include a numerical example.
@@ 1 Introduction and Preliminary Facts
This paper is a continuation and contains some applications of the paper [13]. In [13], we considered the operator generated in $L_{2}[0,1]$ by the expression
$y^{\prime\prime}+q(x)y,$ (1)
and the boundary conditions
$y(1)=y(0)=0,$
where $q(x)$ is a complex-valued summable function. In the present paper, for brevity of some calculation, $L_{2}[0,1]$ is replaced by $L_{2}[0,\pi]$ and it is assumed that ${\displaystyle\int_{0}^{\pi}}q(x)dx=0$. Therefore, we consider the operator $L(q)$ generated in $L_{2}[0,\pi]$ by the expression (1) and the boundary conditions
$y(\pi)=y(0)=0,$
where $q(x)$ is a complex-valued summable function.
In [13], we proved that the eigenvalue $\lambda_{n}$ of the operator $L(q)$ lying in $O(1)$ neighborhood of $n^{2}$ satisfies the equality
$(\lambda_{n}-n^{2}-A_{m}(\lambda_{n}))(\Psi_{n}(x),\sin nx)=R_{m+1},$ (2)
where $n$ is a large number, $\Psi_{n}(x)$ is an eigenfunction corresponding to the eigenvalue $\lambda_{n}$ and satisfying the inequality $|(\Psi_{n},\sin nx)|>1/2$,
@@
and $C_{n}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)\cos nxdx$. Here the sums for $a_{k}(\lambda_{n})$ and $R_{m+1}(\lambda_{n})$ are taken under the conditions
 $\text{ }\sum\limits_{j=1}^{s}n_{j}\neq 0,-2n,$
for s=1,2,,ks=1,2,...,k and s=1,2,,m+1s=1,2,...,m+1, respectively. Moreover, we have
@@
In this paper, in Section 2, we improve the asymptotic formulas, instead of the series in (8), we write some Fourier coefficients of some functions explicitly expressed by the potential $q$. For the potential of bounded variation, we explicitly determine the first and second terms of the asymptotic formulas, which have the order $O(n^{-1})$ and $O(n^{-2})$, respectively (see (33) and Theorem 1). Note that, generally speaking, if the potential $q$ is a piecewise continuous function with some jump, then the first term $C_{2n}$ is of order $n^{-1}$ and the next term in the asymptotic formula (33) is of order $n^{-2}$. Consequently, asymptotic formula (33) with error $O(n^{-3})$ explicitly separates the first and second terms from errors. Then, we demonstrate this formula for the Kronig-Penney potential
@@
where $q(x+\pi)=q(x)$, and $c\in(0,\pi)$. Without loss of generality, we assume that $a<b$, and
$q_{0}=\frac{1}{\pi}\int_{0}^{\pi}q(x)dx=0.$
Therefore, we have
$ac+b(\pi-c)=0,\qquad a<0<b,$ (10)
and $(b-a)c=b\pi$.
Note that, the Kronig-Penney model is a simplified model for an electron in a one-dimensional periodic potential and has been studied in many works (see, for example, [1, 3, 8, 10, 11], and references therein). Veliev [10, 11] studied the bands and gaps in the spectrum of the Schrödinger operator with Kronig-Penney potential and obtained asymptotic formulas for the length of the gaps in the spectrum.
In Section 3, we provide estimates for small Dirichlet eigenvalues of the Schrödinger operator $L(q)$ with the Kronig-Penney potential. We obtain some useful equations for calculating the Dirichlet eigenvalues using Rouche’s theorem. These equations are derived from some iterations formulas by the methods used in [7, 9, 12, 13]. It is important to note that in these papers, the authors used asymptotic formulas for large eigenvalues, which cannot be applied to small eigenvalues. Therefore, in Section 3, we examine the small eigenvalues using numerical methods. Moreover, as a result of meticulous and detailed investigations, we establish conditions on the potential under which all small eigenvalues satisfy the equations obtained from the iterative formulas.
@@ 2 On the Sharp Asymptotic Formulas
$A_{1,n}(n^{2}):=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}^{2}}{n^{2}-(n+k)^{2}}$ (11)
and
$B_{1,n}(n^{2})=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}C_{k+2n}}{n^{2}-(n+k)^{2}}.$ (12)
Then, we consider these series in the following lemmas. First, let us consider $A_{1,n}(n^{2})$. For this introduce the notations:
@@
where
$q_{k}=\frac{1}{\pi}\int_{0}^{\pi}q(x)e^{-ikx}dx\,,\qquad q_{0}=0.$
@@
$D_{1}(n,f)=:\frac{i}{4n\pi}\int_{0}^{\pi}f(x)(Q(x,n)-Q_{n,0})e^{-i2nx}\,dx,\text{ }$
$D_{2}(n,f)=:\frac{i}{4n\pi}\int_{0}^{\pi}f(x)G(x,n)e^{-i2nx}\,dx,$
@@
Since
$\frac{1}{k(2n+k)}=\frac{1}{2n}\biggl(\frac{1}{k}-\frac{1}{2n+k}\biggr),$
we see that
@@
Now let us estimate the first term on the right side of (16). Since $q_{-k}=\overline{q_{k}}$ and $\cos kx=\frac{1}{2}(e^{ikx}+e^{-ikx})$ we have
$C_{k}^{2}=\frac{1}{4}(2\left|q_{k}\right|^{2}+q_{k}^{2}+q_{-k}^{2}).$
Therefore, the first term on the right side of (16) can be written in the form
@@
It is clear that
$Q(\pi,n)=\int_{0}^{\pi}q(t)e^{i2nt}\,dt-q_{-2n}\pi=0,\qquad Q(0,n)=0$
and that the derivative of $Q(x,n)$ with respect to $x$ is
$Q^{\prime}(x,n)=q(x)e^{i2nx}\,-q_{-2n}.$
Therefore, using the integration by parts, we see that
@@
$G(0,n)=0$ and
$G^{\prime}(x,n)=q(x)e^{i2nx}\,+\frac{i\pi}{2}e^{ix}q_{-2n}.$
Therefore,
@@
Thus the proof of (15) follows from (16)-(21). The lemma is proved.
@@
Proof. To prove (22), we estimate the left and write side of this equality separately. First estimate the right-side. Let $Q_{k}={\displaystyle\int_{0}^{\pi}}Q(x)\sin kxdx$. Then, $Q_{k}=\pi C_{k}/k$. Therefore, we have the decomposition
@@
of $Q(x)$. Using the equality $(Q^{2}(x))^{\prime}=2q(x)Q(x)$, the decomposition of $Q(x)$, and integration by parts, and then the equality $\sin kx\sin 2nx=\frac{1}{2}(\cos(2n-k)x-\cos(2n+k)x)$, we obtain
@@
$\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx=\frac{-\pi}{n}\sum_{k\in\mathbb{Z},k\neq 0}\frac{C_{k}C_{2n-k}}{k}$
@@
Now, we prove that the left side of (22) also can be written in this form. Doing the substitution $k\rightarrow-k$, and then using the equality
$\frac{1}{k(2n-k)}=\frac{1}{2n}\biggl(\frac{1}{k}+\frac{1}{2n-k}\biggr),$
we see that
@@
$\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{2n-k}=\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{k}.$
Therefore, we have
$B_{1,n}(n^{2})=\frac{1}{n}\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{k},$
@@
$a_{2,1}(n^{2})=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{k+l}}{[n^{2}-(n+k)^{2}][n^{2}-((n+k+l))^{2}]}$
$\qquad=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(2n+k)(k+l)(2n+k+l)}$
and
@@
Using the equality
$\frac{1}{(k+l)(2n+k+l)}=\frac{1}{2n}\left(\frac{1}{k+l}+\frac{1}{2n+k+l}\right),$
for $l=0$ and $l\neq 0$, we obtain
@@
$S_{1}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(k+l)},$
$S_{2}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{(2n+k)(k+l)},$
$S_{3}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(2n+k+l)},$
and
$S_{4}(n,q)=\sum\limits_{k,k+l\neq 0,2n}\frac{C_{k}C_{l}C_{-k-l}}{(2n+k)(2n+k+l)}.$
@@
with the equal multiplicands,
$C_{k}C_{l}C_{-k-l},\quad C_{l}C_{k}C_{-k-l},\quad C_{k}C_{-k-l}C_{l},$
$C_{l}C_{-k-l}C_{k},\quad C_{-k-l}C_{l}C_{k},\quad C_{-k-l}C_{k}q_{l}$
and using the equality
$\frac{1}{k(k+l)}+\frac{1}{l(k+l)}=\frac{1}{kl},$
we obtain that
@@
$C_{n}=\int_{0}^{\pi}q(x)\cos nxdx=O\bigl(\frac{1}{n}\bigr)$ (26)
as $n\rightarrow\infty$. For example, if $q$ is a function of bounded variation on $0,\pi]$, then (26) holds. There is a larger set of functions satisfying (26) (see [14]). Here we do not discuss the set $E$, since this is not the aim of this paper. We will apply the obtained further result for potential (9) which is a bounded variation.
@@
Now, we estimate $S_{j}(n,q)$, for $j=2,3,4$, in the following lemma. For this, we use the following easily checkable relations
$\sum\limits_{k\neq 0,-2n}\frac{1}{k^{2}(2n+k)}=O\bigl(\frac{1}{n}\bigr),$ (27)
$\sum\limits_{l\neq 0,-k}\frac{1}{l(k+l)^{2}}=O\bigl(\frac{1}{k}\bigr),$ (28)
and
$\sum\limits_{k\neq 0,-2n}\frac{1}{\left|k(2n+k)\right|}=O\bigl(\frac{\ln n}{n}\bigr).$ (29)
To prove (27), split the sum into three parts
@@
and estimate each of them in a standard way. Equalities (28) and (29) can be proved in the same way.
@@
if $q\in E$. Now let us estimate $a_{2,2}(n^{2})$. Arguing as in the proof of (23), we see that
$a_{2,2}(n^{2})=\frac{1}{4n^{2}}(H_{1}+H_{2}+H_{3}+H_{4}),$
where
$H_{1}(q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{k(k+l)},$
$H_{2}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{(2n+k)(k+l)},$
$H_{3}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{k(2n+k+l)},$
and
$H_{4}(n,q)=\sum\limits_{k,k+l\neq 0,2n}\frac{C_{k}C_{l}C_{2n+k+l}}{(2n+k)(2n+k+l)}.$
@@
$\lambda_{n}=n^{2}-C_{2n}-\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx+D(n,q)+2\operatorname{Re}D(n,p)+O\bigl(\frac{1}{n^{3}}\bigr),$ (33)
@@
Proof. It follows from (8), (26) and (27) that
$\lambda_{n}=n^{2}+O\bigl(\frac{1}{n}\bigr).$
Using this, we easily get the estimate
$\sum_{\begin{subarray}{c}k\neq 0,-2n\end{subarray}}^{\infty}\mid\frac{1}{\lambda_{n}-(n+k)^{2}}-\frac{1}{n^{2}-(n+k)^{2}}\mid=O\bigl(\frac{1}{n^{3}}\bigr).$
@@
for $j=1,2,\ldots$. Therefore it follows from (2)-(6) that
@@
where
$a_{1}\left(n^{2}\right)=D(n,q)+2\operatorname{Re}D(n,p)+O\bigl(\frac{1}{n^{3}}\bigr)$ (35)
(see Lemma 1, Lemma 2 and (26)),
@@
Moreover, arguing as in the proof of (36) we see that $a_{3}\left(n^{2}\right)=O\bigl(\dfrac{1}{n^{3}}\bigr)$. Thus, the proof of the theorem follows from (34)-(36).
@@
where $Q(x,n)$, and $G(x,n)$ are defined by (13) and (14). Therefore, by Lemma 1 we have
@@
We first calculate the terms in $D_{1}(n,q)$. Using $(b-a)c=b\pi$ or $ac=b(c-\pi)$ by (10) and $q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$, we have
@@
$=\frac{(b-a)(e^{i2nc}-1)}{4n^{2}}+\pi(\frac{\pi}{2}-c)q_{-2n}=\frac{(b-a)(e^{i2nc}-1)}{4n^{2}}+\pi(\frac{\pi}{2}-c)\frac{i(b-a)(e^{i2nc}-1)}{2\pi n}$
$=\frac{(b-a)(e^{i2nc}-1)}{4n^{2}}-\frac{i\pi(b+a)(e^{i2nc}-1)}{4n},$ (37)
@@
and substituting $q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$ in the last expression, we have
$D_{1}(n,q)=-\frac{ab}{8n^{2}}+\frac{i(b^{2}-a^{2})(e^{-i2nc}-1)}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{16\pi^{2}n^{4}}$
$\qquad+\frac{i(b^{2}-a^{2})(1-\cos 2nc)}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{16\pi^{2}n^{4}}$
$\qquad=-\frac{ab}{8n^{2}}+\frac{(b^{2}-a^{2})\sin 2nc}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{8\pi^{2}n^{4}}.$ (38)
@@
Thus, by (38), (39), and Lemma 1, we write
$D(n,q)=D_{1}(n,q)+D_{2}(n,q)=-\frac{ab}{4n^{2}}+\frac{(b^{2}-a^{2})\sin 2nc}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{8\pi^{2}n^{4}}$
$\qquad+\frac{i(b^{2}-a^{2})(e^{-i2nc}-1)}{16\pi n^{3}}+\frac{ib(b-a)(\cos 2nc-1)}{16n^{3}}-\frac{ib(b-a)(1-e^{-i2nc})(e^{i2nc}+e^{ic})}{16n^{2}(2n-1)}$
$\qquad+\frac{(b-a)^{2}(\cos 2nc-1)}{32\pi n^{4}}+\frac{(b-a)(e^{-i2nc}-1)[(b+a)e^{i2nc}+(b-a)e^{ic}]}{16\pi n^{2}(2n-1)^{2}}$
or
@@
Now, let us calculate $\operatorname{Re}D(n,p)$, where $p(x)$ is defined in Lemma 1. For the Kronig-Penney potential, $p(x)$ becomes
$p(x)=\begin{cases}b&\text{if }x\in[0,\pi-c),\\[5.69054pt] a&\text{if }x\in[\pi-c,\pi]\end{cases}.$
Therefore,

### v2

@@ Abstract
In this paper, we first improve some asymptotic formulas previously obtained and provide sharp asymptotic formulas explicitly expressed by the potential. For the potentials of bounded variation, we obtain asymptotic formulas in which the first and second terms are explicitly determined and separated from the error terms. In addition, we illustrate these formulas for the Kronig-Penney potential. We then provide estimates for the small Dirichlet eigenvalues of the one-dimensional Schrödinger operator in the Kronig-Penney model. We derive several useful equations from certain iteration formulas for computing these Dirichlet eigenvalues, and prove that all the eigenvalues can be found by the fixed point iteration. Then, using the Banach fixed point theorem, we estimate the eigenvalues numerically. Moreover, we present error estimates and include a numerical example.
@@ 1 Introduction and Preliminary Facts
In the present paper, we consider the operator generated in $L_{2}[0,\pi]$ by the expression
$y^{\prime\prime}+q(x)y$ (1)
and the boundary conditions $y(\pi)=y(0)=0$, where $q(x)$ is a complex-valued summable function. In [13], we proved that the eigenvalue $\lambda_{n}$ of the operator $L(q)$ lying in $O(1)$ neighborhood of $n^{2}$ satisfies the equality
$(\lambda_{n}-n^{2}-A_{m}(\lambda_{n}))(\Psi_{n},\sin nx)=R_{m+1},$ (2)
where $n$ is a large number, $\Psi_{n}$ is an eigenfunction corresponding to the eigenvalue $\lambda_{n}$ and satisfying the inequality $|(\Psi_{n},\sin nx)|>1/2$,
@@
$C_{n}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)\cos nxdx$, and, without loss of generality, it is assumed that $C_{0}=0$. Here the sums for $a_{k}(\lambda_{n})$ and $R_{m+1}(\lambda_{n})$ are taken under the conditions $n_{1}+n_{2}+...+n_{s}\neq 0,-2n$, for s=1,2,,ks=1,2,...,k and s=1,2,,m+1s=1,2,...,m+1, respectively. Moreover, we have
@@
In this paper, in Section 2, we improve the asymptotic formulas, instead of the series in (8), we write some Fourier coefficients of some functions explicitly expressed by the potential $q$. For the potential of bounded variation, we explicitly determine the first and second terms of the asymptotic formulas, which have the orders $O(n^{-1})$ and $O(n^{-2})$, respectively (see (30) and Theorem 1). Note that, generally speaking, if the potential $q$ is a piecewise continuous function with some jump, then the first term $C_{2n}$ is of order $n^{-1}$ and the next term in asymptotic formula (30) is of order $n^{-2}$. Consequently, asymptotic formula (30) with error $O(n^{-3})$ explicitly separates the first and second terms from the error terms. Then, we demonstrate this formula for the Kronig-Penney potential
@@
where $q(x+\pi)=q(x)$, and $c\in(0,\pi)$. Without loss of generality, we assume that $a<b$. This inequality with the condition $C_{0}=0$ implies that
$ac+b(\pi-c)=0,\qquad a<0<b,\text{ }(b-a)c=b\pi.$ (10)
Note that, the Kronig-Penney model is a simplified model for an electron in a one-dimensional periodic potential and has been studied in many works (see, for example, [1, 3, 8, 10, 11], and references therein). Veliev [10, 11] studied the bands and gaps in the spectrum of the Schrödinger operator with the Kronig-Penney potential and obtained asymptotic formulas for the length of the gaps in the spectrum.
In Section 3, we provide estimates for small Dirichlet eigenvalues of the Schrödinger operator $L(q)$ with the Kronig-Penney potential. First, we derive equation (42) for calculating the Dirichlet eigenvalues using some iterations formulas by the methods used in [7, 9, 12, 13]. It is important to note that, the main difficulty of this section is to prove that all the eigenvalues (small and large) of $L(q)$ can be obtained from (42) by the fixed point iteration. Then, we examine the small eigenvalues using numerical methods. Moreover, in Section 3, as a result of meticulous and detailed investigations, we establish conditions on the potential under which all the small eigenvalues satisfy equality (42) and the Banach fixed point theorem can be used for this equality.
@@ 2 On the Sharp Asymptotic Formulas
$A_{1,n}(n^{2}):=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}^{2}}{n^{2}-(n+k)^{2}},\qquad B_{1,n}(n^{2}):=\sum_{\begin{subarray}{c}k=-\infty\\ k\neq 0,-2n\end{subarray}}^{\infty}\frac{C_{k}C_{k+2n}}{n^{2}-(n+k)^{2}}.$ (11)
Then, we consider these series in the following lemmas. First, let us consider $A_{1,n}(n^{2})$. For this, introduce the notations:
@@
where $q_{k}=\dfrac{1}{\pi}{\displaystyle\int_{0}^{\pi}}q(x)e^{-ikx}dx$ and $q_{0}=0$.
@@
$D_{1}(n,f)=\frac{i}{4n\pi}\int_{0}^{\pi}f(x)(Q(x,n)-Q_{n,0})e^{-i2nx}\,dx,\quad D_{2}(n,f)=\frac{i}{4n\pi}\int_{0}^{\pi}f(x)G(x,n)e^{-i2nx}\,dx,$
@@
Since $\dfrac{1}{k(2n+k)}=\dfrac{1}{2n}\bigl(\dfrac{1}{k}-\dfrac{1}{2n+k}\bigr)$, we see that
@@
Now let us estimate the first term on the right side of (15). Since $q_{-k}=\overline{q_{k}}$ and $\cos kx=\dfrac{1}{2}(e^{ikx}+e^{-ikx})$, we have $C_{k}^{2}=\dfrac{1}{4}(2\left|q_{k}\right|^{2}+q_{k}^{2}+q_{-k}^{2})$. Therefore, the first term on the right side of (15) can be written in the form
@@
It is clear that $Q(\pi,n)=0$, $Q(0,n)=0$, and that the derivative of $Q(x,n)$ with respect to $x$ is $Q^{\prime}(x,n)=q(x)e^{i2nx}\,-q_{-2n}$. Using the integration by parts, we see that
@@
$G(0,n)=0$ and that $G^{\prime}(x,n)=q(x)e^{i2nx}\,+\dfrac{i\pi}{2}e^{ix}q_{-2n}$. Therefore,
@@
Thus, the proof of (14) follows from (15)-(20). The lemma is proved.
@@
Proof. To prove (21), we estimate the left and right sides of this equality separately. First, estimate the right-side. Let $Q_{k}={\displaystyle\int_{0}^{\pi}}Q(x)\sin kxdx$. Then, $Q_{k}=\pi C_{k}/k$. Therefore, we have the decomposition
@@
of $Q(x)$. Using the equality $(Q^{2}(x))^{\prime}=2q(x)Q(x)$, integration by parts, the decomposition of $Q(x)$, and then the equality $\sin kx\sin 2nx=\dfrac{1}{2}(\cos(2n-k)x-\cos(2n+k)x)$, we obtain
@@
$\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx=\frac{-\pi}{n}\sum_{k\in\mathbb{Z},k\neq 0}\frac{C_{k}C_{2n-k}}{k}.$
@@
Now, we prove that the left side of (21) also can be written in this form. Doing the substitution $k\rightarrow-k$, and then using the equality $\dfrac{1}{k(2n-k)}=\dfrac{1}{2n}\bigl(\dfrac{1}{k}+\dfrac{1}{2n-k}\bigr)$, we see that
@@
$\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{2n-k}=\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{k},\qquad B_{1,n}(n^{2})=\frac{1}{n}\sum_{k\in\mathbb{Z},k\neq 0,2n}\frac{C_{k}C_{2n-k}}{k},$
@@
$a_{2,1}(n^{2})$ $=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{k+l}}{[n^{2}-(n+k)^{2}][n^{2}-((n+k+l))^{2}]}=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(2n+k)(k+l)(2n+k+l)},$
@@
Using the equality $\dfrac{1}{(k+l)(2n+k+l)}=\dfrac{1}{2n}\bigl(\dfrac{1}{k+l}+\dfrac{1}{2n+k+l}\bigr)$, for $l=0$ and $l\neq 0$, we obtain
@@
$S_{1}(n,q)$ $=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(k+l)},\qquad S_{2}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{(2n+k)(k+l)},$
$S_{3}(n,q)$ $=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{-k-l}}{k(2n+k+l)},\qquad S_{4}(n,q)=\sum\limits_{k,k+l\neq 0,2n}\frac{C_{k}C_{l}C_{-k-l}}{(2n+k)(2n+k+l)}.$
@@
with the equal multiplicands, $C_{k}C_{l}C_{-k-l}$, $C_{l}C_{k}C_{-k-l}$, $C_{k}C_{-k-l}C_{l}$, $C_{l}C_{-k-l}C_{k}$, $C_{-k-l}C_{l}C_{k}$, $C_{-k-l}C_{k}C_{l}$ and using the equality $\dfrac{1}{k(k+l)}+\dfrac{1}{l(k+l)}=\dfrac{1}{kl}$, we obtain that
@@
$C_{n}=\frac{1}{\pi}\int_{0}^{\pi}q(x)\cos nxdx=O\bigl(\frac{1}{n}\bigr),$ (25)
as $n\rightarrow\infty$. For example, if $q$ is a function of bounded variation on $0,\pi]$, then (25) holds. There is a larger set of functions satisfying (25) (see [14]). Here we do not discuss the set $E$, since this is not the aim of this paper. We will apply the obtained further results for potential (9) which is a bounded variation.
@@
Now, we estimate $S_{j}(n,q)$, for $j=2,3,4$, in the following lemma. For this, we use the following easily checkable relations:
$\sum\limits_{k\neq 0,-2n}\frac{1}{k^{2}(2n+k)}=O\bigl(\frac{1}{n}\bigr),\quad\sum\limits_{l\neq 0,-k}\frac{1}{l(k+l)^{2}}=O\bigl(\frac{1}{k}\bigr),\quad\sum\limits_{k\neq 0,-2n}\frac{1}{\left|k(2n+k)\right|}=O\bigl(\frac{\ln n}{n}\bigr).$ (26)
To prove the first equality, split the sum into three parts
@@
and estimate each of them in a standard way. The second and third equalities can be proved in the same way.
@@
if $q\in E$. Now, let us estimate $a_{2,2}(n^{2})$. Arguing as in the proof of (22), we see that $a_{2,2}(n^{2})=\dfrac{1}{4n^{2}}(H_{1}+H_{2}+H_{3}+H_{4})$, where
$H_{1}(n,q)$ $=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{k(k+l)},\qquad H_{2}(n,q)=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{(2n+k)(k+l)},$
$H_{3}(n,q)$ $=\sum\limits_{k,k+l\neq 0,-2n}\frac{C_{k}C_{l}C_{2n+k+l}}{k(2n+k+l)},\qquad H_{4}(n,q)=\sum\limits_{k,k+l\neq 0,2n}\frac{C_{k}C_{l}C_{2n+k+l}}{(2n+k)(2n+k+l)}.$
@@
$\lambda_{n}=n^{2}-C_{2n}+\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx+D(n,q)+2\operatorname{Re}D(n,p)+O\bigl(\frac{1}{n^{3}}\bigr),$ (30)
@@
Proof. It follows from (8), (25) and (26) that $\lambda_{n}=n^{2}+O\bigl(\dfrac{1}{n}\bigr)$. Using this, we easily get the estimate
$\sum_{\begin{subarray}{c}k\neq 0,-2n\end{subarray}}^{\infty}\biggl|\frac{1}{\lambda_{n}-(n+k)^{2}}-\frac{1}{n^{2}-(n+k)^{2}}\biggr|=O\bigl(\frac{1}{n^{3}}\bigr).$
@@
for $j=1,2,\ldots$. Therefore, it follows from (2)-(6) that
@@
where $a_{1}(n^{2})=A_{1,n}(n^{2})-B_{1,n}(n^{2})$, equivalently
$a_{1}\left(n^{2}\right)=\frac{1}{\pi}\int_{0}^{\pi}Q^{2}(x)\cos 2nxdx+D(n,q)+2\operatorname{Re}D(n,p)+O\bigl(\frac{1}{n^{3}}\bigr)$ (32)
(see Lemma 1, Lemma 2 and (25)), and
@@
Moreover, arguing as in the proof of (33), we see that $a_{3}\left(n^{2}\right)=O\bigl(\dfrac{1}{n^{3}}\bigr)$. Thus, the proof of the theorem follows from (31)-(33).
@@
where $Q(x,n)$, and $G(x,n)$ are defined by (12) and (13). Therefore, by Lemma 1, we have
@@
We first calculate the terms in $D_{1}(n,q)$. Using $(b-a)c=b\pi$ or $ac=b(c-\pi)$ by (10) and substituting $q_{-2n}=\dfrac{i(b-a)(e^{i2nc}-1)}{2\pi n}$ by (34), we have
@@
$=\frac{(b-a)(e^{i2nc}-1)}{4n^{2}}+\pi(\frac{\pi}{2}-c)q_{-2n}=\frac{(b-a)(e^{i2nc}-1)}{4n^{2}}-\frac{i\pi(b+a)(e^{i2nc}-1)}{4n},$
@@
and by (34), we have
$D_{1}(n,q)=-\frac{ab}{8n^{2}}+\frac{(b^{2}-a^{2})\sin 2nc}{16\pi n^{3}}-\frac{(b-a)^{2}(1-\cos 2nc)}{8\pi^{2}n^{4}}.$ (35)
@@
Therefore, using (35), (36), Lemma 1, and simplifying a bit, we obtain
@@
Now, let us calculate $\operatorname{Re}D(n,p)$, where $p(x)$ is defined in Lemma 1. For the Kronig-Penney potential, $p(x)$ becomes $p(x)=b$ for $x\in[0,\pi-c)$ and $p(x)=a$ for $x\in[\pi-c,\pi]$. Therefore, we have